l-UNCOUPLING AND ROTATIONAL STRUCTURE IN THE HIGH-RYDBERG STATES OF HCO \underline{HARTMUT G. HEDDERICH}, ERIC E. MAYER and EDWARD R. GRANT, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

Rotational structure associated with autoionizing states of HCO in the n = 13 complex is analyzed for core vibrational states (010) and (001). l-uncoupling is observed.